OpenStudy (anonymous):

One litre of gasoline releases 35 MJ of energy when burned. The automobile's efficiency is about 40%. If the speed of the car is 9.0m*s^-1 when the engine power is at 20 kW, calculate the distance the car can travel on one litre of gasoline at this speed.

OpenStudy (shane_b):

I have to start off by assuming "engine power is at 20kW" is talking about engine OUTPUT power. If that's supposed to be the engine's input power, then you'll have to adjust my calcs a bit. First, calculate the input power if the output power of the engine is 20kW and it's only 40% efficient:\[E_{ff}=\frac{P_{out}}{P_{in}}=\frac{20kW}{P_{in}}=0.4\]\[P_{in}=50kW=50kJ/s\] Now let's figure out how many seconds the engine can run when it's consuming energy at a rate of 50kJ/s:\[\frac{35MJ}{50k J/s}=700s\]Since you can only travel for 700s on 1L of fuel at a speed of 9m/s, you end up with a distance of:\[700s*9m/s=6300m=6.3km\]